# Hug Reports: Supporting Expression of Appreciation between Users and Contributors of Open Source Software Packages

### v1

@@ 2.4. Lowering effort in expressing appreciation while preserving meaningfulness
Central to our approach is the idea of lowering the effort for users to express appreciation, by fine-tuning the system to the manner in which software is used in the development environment. Before we can proceed, however, we must address an apparent dilemma: while lowering the effort involved in expressing appreciation may encourage appreciation, it can also undermine its meaningfulness. While low effort actions can encourage interactions, prior work suggests these interactions can feel limited in the authenticity (Monroy-Hernández et al., 2011; Zhang et al., 2022a) and support (Wohn et al., 2016) they convey. Is lowering effort bound to dilute the meaningfulness of appreciation?
@@
In our work, it is specifically the procedural effort that we attempt to lower. Our approach attempts to lower procedural effort by locating a communication affordance within the code editor and by allowing users to express appreciation at the level of packages, modules, or functions rather than individual commits or pull-requests. At the same time, we give users control over the amount of personal effort they invest by allowing them to customize their appreciation messages.
@@ 3.1. Concept proposal
In envisioning solutions that would overcome barriers in expressing appreciation, we arrived at a conceptual design proposal for the Hug Reports [^9] We chose the name to suggest to suggest an inversion of the concept of bug reports in software development, which are intended to convey critical feedback. system, a unidirectional communication system that would afford users the ability to select a package currently in use and direct a “thanks” towards contributors from within their code editors. The affordance would allow users to direct thanks towards entire packages (e.g. matplotlib), their modules (e.g. pyplot), or specific functions within those modules (e.g. pyplot.scatter). To map the thanked packages to contributors, we planned to use activity traces from GitHub repositories corresponding to the packages to identify the relevant contributors. Finally, we would deliver the “thanks” to contributors on behalf of the users.
The key decisions in Hug Reports were to: (1) lower procedural effort by locating a communication affordance within the code editor and allowing users to express appreciation at the level of packages, modules, or functions rather than individual commits or pull-requests; and (2) provide users a subtle reminder of the contributors through the material presence of the affordance in the code editor. As we describe in Section 2.3, these decisions differentiate our approach from prior appreciation systems, and especially those in open source. So, we wanted to investigate the merits of these key decisions. At the same time, due to our departure from previous approaches, we anticipated that unexpected cross-cutting constraints and requirements would reveal themselves when an end-to-end system is deployed.
@@ 3.2. Overview of method and research questions
At this early stage of the design process, then, we required a method to assess the feasibility of the approach and to identify cross-cutting concerns. We wanted to rapidly explore the design space and iterate on our design concept, while reducing the risk of developing an end-to-end technical system that users and contributors did not ultimately want So, we chose to use the method of technology probes. As per Hutchinson et al.  (Hutchinson et al., 2003), technology probes are functioning technological artifacts that balance three goals: social science: “understanding the use and the users”; engineering: “field testing the technology”; and design: “inspiring users to think of new kinds of technology to support their needs”. Technology probes are often used in the design of social applications, to engage participants early in the design process (Jörke et al., 2023; Sellen et al., 2006; Leong et al., 2023; Khadpe et al., 2024) and to find out about the “unknown” when deployed (Hutchinson et al., 2003). So, we chose to develop a barebones version of Hug Reports as a technology probe and deploy it in a field study to address the following research questions:
  • RQ1: How do the key decisions of Hug Reports impact expression of appreciation?
    • RQ1-U1: To what extent does lowering procedural effort—locating a communication affordance within the code editor and allowing users to express appreciation at the level of packages, modules, or functions—encourage users to express appreciation?
    • RQ1-U2: To what extent does subtly reminding users of the contributors, from within the code editor, encourage users to express appreciation?
    • RQ1-C: To what extent do contributors find the appreciation meaningful?
    These questions revolve around the probe’s goal of “field testing” the key decisions.
  • RQ2: How is appreciation felt, expressed, received, and interpreted?
@@
    • RQ2-U: How did users express appreciation, in terms of when they felt appreciation, what they felt appreciation towards, and how they expressed it?
    • RQ2-C: How were contributors’ perceptions of the appreciation affected by what was being appreciated, how appreciation was expressed, and the broader culture of open source?
    These questions revolve around the probe’s goal of “understanding the use and users”. They allow us to take an end-to-end view of the system and identify cross-cutting concerns and constraints.
  • RQ3: What opportunities do users and contributors envision for improving appreciation, given their respective practices of using and contributing?
@@
    This question revolves around the probe’s goal of “inspiring users to think of new kinds of technology to support their needs”.
@@
In our research activities, we planned to involve two populations of participants: users of open source software packages and contributors to open source software packages. Hereafter, we slightly overload terms and refer to participants whom we planned to involve/were involved in their capacity as users, simply as “users”. Similarly, we refer to participants whom we planned to involve/were involved in their capacity as contributors to packages, simply as “contributors”. For the purpose of the study, we decided to develop realistic versions of the two participant-facing components of the concept: an extension for code editors through which users would express appreciation and the notification system that would convey the appreciation to contributors.
@@ 3.3. Considerations in implementing the probe and designing the study
(2) Minimizing disruption to contributors. Following from the above, since contributors would only be introduced to the study by the actual notifications they received, we decided to deliver these privately and individually via email. We wanted to avoid any reputational consequences that might result from notifying them through public channels. Further, to reduce interruptions to contributors from our study, we decided to contact them with only one notification. Synchronously sending every “thanks” to contributors could be disruptive if there were multiple messages directed to them. As a result, we decided to aggregate the “thanks” expressed by users across a period of time before notifying contributors. This decision resulted in a sequencing of research activities: first, a deployment of the code editor extension with users, followed by notifying the contributors. Finally, we decided to limit the notifications to recent contributors so as to not disturb past contributors who might have disengaged from the project. As a starting point, we decided to send notifications to only the 20 most recent contributors of the thanked software, if there were more than 20 unique contributors. Both users and contributors were made aware of this heuristic throughout the study. The decision to send the thanks after aggregation, and the decision to notify the 20 most recent contributors were simple options that met our study requirements, and allowed us to deploy a working probe to investigate our research questions. We do not intend to suggest that these are the best choices for a final appreciation system (see our note on implementation details in Appendix A).
As a result of these considerations, we decided to pursue our probe development and study activities sequentially: first, we deployed the code editor extension with users, and then we aggregated the thanks messages and notified the contributors.
@@ 4.1. The Hug Reports extension
We developed an extension for the Visual Studio Code (VS Code) editor that allows users to express appreciation to developers of Python and JavaScript/TypeScript packages that they are using [^10] JavaScript and Python are the most commonly used programming languages according to the 2023 Stack Overflow Developer Survey (Overflow, 2023). Similarly, we chose to develop an extension for Visual Studio Code because it is the most popular development environment (Overflow, 2023).. It also allows them to specify whether the object of their appreciation is the entire package, specific modules, or functions within the package. The extension gets activated whenever a Python (.py), JavaScript (.js, .jsx), or TypeScript (.ts, .tsx) file is opened. The user flow of the extension is described in Figure 2. The main communication affordance of the extension is a button ([Uncaptioned image]) that is rendered in the gutter, on every line of the file that interfaces with an imported package. The button is present next to imports of entire packages (e.g. “import Quill from "quill";”), imports of specific modules (e.g. “from matplotlib import pyplot as plt”), as well as lines where a function from the package or module is being used (e.g. “img = cv2.imread(’watch.jpg’,cv2.IMREAD_GRAYSCALE)”). When a user right-clicks the button on a given line, it displays a contextual menu with an option to “Say Thanks”. Clicking that option logs a thanks along with the line of code, which is used as a representation of the object of the users appreciation. Clicks next to the import of a package are treated as thanks directed at the package as a whole. Similarly clicks next to imports of modules, and next to function calls, are treated as being directed at the module and function respectively. The thanks itself, is a one-bit signal of appreciation, similar to a “like” or “upvote”. A modal pop-up notifies the user that their thanks has been logged and gives them the option to “Say More”. Clicking “Say More” redirects them to a web form in their browser where they can type out a longer personal note. Users are not required to create an account. The thanks as well as personal notes do not identify their author; they are only associated with an installation ID, which is uniquely assigned for every installation of the extension. Each thanks is logged to a database and is associated with the installation ID of the user, the line number, the line of code, and a personal note if provided. The implementation of the extension is open source and available at: https://github.com/Hug-Reports/hug-reports-extension-v0. Implementation notes are in Appendix A.
@@
By providing a communication affordance within the code editor, and allowing users to express appreciation at the level of packages, modules, or functions rather than individual commits or pull-requests, the extension lowers procedural effort in expressing appreciation. We chose to present the communication affordance via a persistent button on the interface to provide users a subtle reminder of the contributors.
@@ 4.2.1. Participants
Two criteria were used to screen potential participants. We required that they primarily used VS Code as their development environment and that they anticipated writing Python/JavaScript/TypeScript code at least 3 times a week. The screening survey also included questions about users’ demographics and backgrounds. Responses are summarized in Table 1. The 18 users (12 men and 6 women) were aged 23-30. The survey asked users to report their experience levels as either Beginner, Advanced Beginner, Intermediate, Advanced, or Expert (categories present in the developer skill matrix). As Table 1 shows, users described themselves as Advanced (13/28), Intermediate (4/18), and Expert (1/18). We also asked users to report their weekly average of the amount of time they spent writing code, which ranged from 5 hours to 45 hours a week.
@@ 4.2.2. Procedure
We deployed the extension for three weeks so that users would have sufficient time to familiarize themselves with it and explore how they might interact with it in the course of their usual programming activities. To observe naturalistic use, required usage was deliberately kept minimal; however, we still encouraged (without enforcing) users to engage with the extension a meaningful amount so that they had opportunities to reveal their experiences, and so that there would be sufficient messages with which to understand contributor experiences. Each user was required to participate in a 15-minute onboarding session conducted over video call, during which we helped them install the extension and provided a brief tutorial. During this session, they were also required to complete a brief pre-study questionnaire that asked them to provide: (1) an open-ended response describing how often they typically thanked contributors of packages they used; and (2) a rating on a seven point scale (strongly disagree to strongly agree) indicating their agreement with the statement: “There are many developers whose work I am grateful for.” These questions were intended to capture their current feelings and practices of expressing appreciation. These are presented in in Table 3 (Appendix B) to provide more detail about the participant population. At the end of the onboarding session, we encouraged users to send thanks at least two times every day they found themselves coding, while expressing that this would not be enforced. During the three weeks, for every thanks, the extension logged the installation ID [^11] For the purpose of the study, we established a link between participants and their installation IDs during the onboarding session., line of code, line number, and personal note if added. Users were compensated \$20 for participating in the onboarding session and completing the pre-study questionnaire. At the end of the three weeks, we invited all 18 users to participate in a one-on-one interview, 14 of whom agreed to participate (those who did not are indicated in Table 1). Interviews were conducted in English, remotely via Zoom, and lasted between 20 and 50 minutes, for which users were compensated \$15. Interviews were semi-structured and investigated factors that contributed to users’ feelings of appreciation and decisions to express it, the experience of selecting what to thank and what to include in personal notes, trends in their usage, and feedback on the interface and interaction. Across these topics, we asked users if there were ways in which different or new designs could have better supported their experience. All interviews were recorded and transcribed.
We planned to use data from the interviews with users to address RQ1-U1, RQ1-U2, RQ2-U, and RQ3. To address RQ2-U, we also planned to analyze usage data. Finally, we planned to use the thanks messages created over the course of the deployment to pursue our research activities with contributors.
@@ 4.3.1. Identifying contributors:
We first identified each unique object that was thanked (content of the line of code at which the button was clicked). For this process, we treated packages, modules, and functions as separate objects since the 20 most recent contributors to each could be different. Thanks logged next to the import of a package were treated as thanks directed at the package as a whole. Similarly thanks logged next to imports of modules, and next to function calls, are treated as being directed at the module and function respectively. We note that each object could have been thanked multiple times (e.g. “import cv2” was thanked twice). Corresponding to each unique object, we maintained a count of the number of thanks it was associated with and all personal notes associated with that object. We identified a total of $7070$ unique objects that were thanked, most just once. We then mapped each object of appreciation to its source code on GitHub. Thanked packages were mapped to their repositories, while thanked modules and functions were mapped to their corresponding file. We used this to then find the 20 most recent contributors. For thanked packages, we identified the 20 most recent contributors to the entire repository. For thanked modules and functions, we identified the 20 most recent contributors to the corresponding file. Every commit, contains an email address of the author of the code patch, which provided us with contact information for the contributors. In determining the 20 most recent contributors, we skipped commits where the provided email address was anonymized (emails ending in “users.noreply.github.com”). If the total number of unique contributors was fewer than 20, we recorded all contributors. This process resulted in a total of $550550$ contributors to be notified. $470470$ had been thanked for $11$ object each, with the remaining being thanked for $22$ to $88$ objects each.
@@ 4.3.2. Notification
As shown in Figure 3 (B), the notification was organized so that each segment showed one object that the contributor had been thanked for (B1). Above the object, the notification showed the number of thanks corresponding to it (B3). Below the object, we included any personal notes associated with it (B4). The notification was delivered via an email (as shown in Figure 3 (A)) that provided the contributors with context about our project, so that they could understand the conditions under which the users had directed these thanks and personal notes towards them. The design of the notification message was modeled after notifications in Wikipedia’s “Thanks” system (Matias et al., [n. d.]).
Figure 3. (A) The Hug Report notification was delivered via an email that provided the contributors with context about our project, so that they could understand the conditions under which the users had directed these thanks and personal notes towards them. (B) The Hug Report notification, itself, was organized so that each segment (B1) showed one object that the contributor had been thanked for (B2). Above the object, the notification showed the number of thanks corresponding to it (B3). Below the object, were any personal notes associated with it (B4)

### v2

@@ 2.4. Lowering effort in expressing appreciation while preserving meaningfulness
Central to our approach is the idea of lowering the effort for users to express appreciation, by fine-tuning the system to the manner in which software is used in the development environment. Before we can proceed, however, we must address an apparent dilemma: while lowering the effort involved in expressing appreciation may encourage appreciation, it can also undermine its meaningfulness. While low-effort actions can encourage interactions, prior work suggests these interactions can feel limited in the authenticity (Monroy-Hernández et al., 2011; Zhang et al., 2022a) and support (Wohn et al., 2016) they convey. Is lowering effort bound to dilute the meaningfulness of appreciation?
@@
In our work, it is specifically the procedural effort that we attempt to lower. Our approach attempts to lower procedural effort by locating a communication affordance within the code editor and by allowing users to express appreciation at the level of packages, modules, or functions rather than individual commits or pull requests. At the same time, we give users control over the amount of personal effort they invest by allowing them to customize their appreciation messages.
@@ 3.1. Concept proposal
In envisioning solutions that would overcome barriers in expressing appreciation, we arrived at a conceptual design proposal for the Hug Reports [^9] We chose the name to suggest an inversion of the concept of bug reports in software development, which are intended to convey critical feedback. system, a unidirectional communication system that would afford users the ability to select a package currently in use and direct a “thanks” towards contributors from within their code editors. The affordance would allow users to direct thanks towards entire packages (e.g. matplotlib), their modules (e.g. pyplot), or specific functions within those modules (e.g. pyplot.scatter). To map the thanked packages to contributors, we planned to use activity traces from GitHub repositories corresponding to the packages to identify the relevant contributors. Finally, we would deliver the “thanks” to contributors on behalf of the users.
The key decisions in Hug Reports were to: (1) lower procedural effort by locating a communication affordance within the code editor and allowing users to express appreciation at the level of packages, modules, or functions rather than individual commits or pull-requests; and (2) provide users a subtle reminder of the contributors through the material presence of the affordance in the code editor. As we describe in Section 2.3, these decisions differentiate our approach from prior appreciation systems, especially those in open source. So, we wanted to investigate the merits of these key decisions. At the same time, due to our departure from previous approaches, we anticipated that unexpected cross-cutting constraints and requirements would reveal themselves when an end-to-end system is deployed.
@@ 3.2. Overview of method and research questions
At this early stage of the design process, then, we required a method to assess the feasibility of the approach and to identify cross-cutting concerns. We wanted to rapidly explore the design space and iterate on our design concept while reducing the risk of developing an end-to-end technical system that users and contributors did not ultimately want So, we chose to use the method of technology probes. As per Hutchinson et al.  (Hutchinson et al., 2003), technology probes are functioning technological artifacts that balance three goals: social science: “understanding the use and the users”; engineering: “field testing the technology”; and design: “inspiring users to think of new kinds of technology to support their needs”. Technology probes are often used in the design of social applications, to engage participants early in the design process (Jörke et al., 2023; Sellen et al., 2006; Leong et al., 2023; Khadpe et al., 2024) and to find out about the “unknown” when deployed (Hutchinson et al., 2003). So, we chose to develop a barebones version of Hug Reports as a technology probe and deploy it in a field study to address the following research questions:
  • RQ1How do the key decisions of Hug Reports impact expression of appreciation? To what extent does lowering procedural effort encourage users to express appreciation? To what extent does reminding users of the contributors encourage users to express appreciation? To what extent do contributors find the appreciation meaningful? (This RQ revolves around the probe’s goal of “field testing” the key decisions.)
@@
  • RQ2How is appreciation felt, expressed, received, and interpreted? How did users express appreciation, in terms of when they felt appreciation, what they felt appreciation towards, and how they articulated it? How were contributors’ perceptions of the appreciation affected by what was being appreciated and how appreciation was articulated? (This RQ revolves around the probe’s goal of “understanding the use and users”.)
@@
  • RQ3What opportunities do users and contributors envision for supporting appreciation? (This RQ revolves around the probe’s goal of “inspiring users to think of new kinds of technology to support their needs”.)
@@
In our research activities, we planned to involve two populations of participants: users of open source software packages and contributors to open source software packages. Hereafter, we slightly overload terms and refer to participants whom we planned to involve/were involved in their capacity as users, simply as “users”. Similarly, we refer to participants whom we planned to involve/were involved in their capacity as contributors to packages, simply as “contributors”. For the study, we decided to develop realistic versions of the two participant-facing components of the concept: an extension for code editors through which users would express appreciation and the notification system that would convey the appreciation to contributors.
@@ 3.3. Considerations in implementing the probe and designing the study
(2) Minimizing disruption to contributors. Since contributors would only learn about the study through the notifications they received, we decided to deliver these privately via individual emails to avoid potential reputational effects of public notifications. To minimize disruption, we sent only one aggregated notification to each contributor with “thanks” across a period of time, rather than notifying them each time a “thanks” was received. This decision resulted in a sequencing of research activities: first, a deployment of the code editor extension with users, followed by notifying the contributors. Finally, we decided to limit the notifications to recent contributors so as to not disturb past contributors who might have disengaged from the project. As a starting point, we decided to send notifications to only the 20 most recent contributors of the thanked software, if there were more than 20 unique contributors. Both users and contributors were made aware of this heuristic throughout the study. The decision to send the thanks after aggregation and the decision to notify the 20 most recent contributors were simple options that met our study requirements and allowed us to deploy a working probe to investigate our research questions. We do not intend to suggest that these are the best choices for a final appreciation system (see our note on implementation details in Appendix A).
As a result of these considerations, we decided to pursue our probe development and study activities sequentially: first, we deployed the editor extension with users, and then we aggregated the thanks messages and notified contributors.
@@ 4.1. The Hug Reports extension
We developed an extension for the Visual Studio Code (VS Code) editor that allows users to express appreciation to developers of Python and JavaScript/TypeScript packages that they are using [^10] JavaScript and Python are the most commonly used programming languages according to the 2023 Stack Overflow Developer Survey (Overflow, 2023). Similarly, we chose to develop an extension for Visual Studio Code because it is the most popular development environment (Overflow, 2023).. It also allows them to specify whether the object of their appreciation is the entire package, specific modules, or functions within the package. The extension gets activated whenever a Python (.py), JavaScript (.js, .jsx), or TypeScript (.ts, .tsx) file is opened. The user flow of the extension is described in Figure 2. The main communication affordance of the extension is a button ([Uncaptioned image]) that is rendered in the gutter, on every line of the file that interfaces with an imported package. The button is present next to imports of entire packages (e.g. “import Quill from "quill";”), imports of specific modules (e.g. “from matplotlib import pyplot as plt”), as well as lines where a function from the package or module is being used (e.g. “img = cv2.imread(’watch.jpg’,cv2.IMREAD_GRAYSCALE)”). When a user right-clicks the button on a given line, it displays a contextual menu with an option to “Say Thanks”. Clicking that option logs a thanks along with the line of code, which is used as a representation of the object of the user’s appreciation. Clicks next to the import of a package are treated as thanks directed at the package as a whole. Similarly clicks next to imports of modules, and next to function calls, are treated as being directed at the module and function respectively. The thanks itself, is a one-bit signal of appreciation, similar to a “like” or “upvote”. A modal pop-up notifies the user that their thanks has been logged and gives them the option to “Say More”. Clicking “Say More” redirects them to a web form in their browser where they can type out a longer personal note. Users are not required to create an account. The thanks as well as personal notes do not identify their author; they are only associated with an installation ID, which is uniquely assigned for every installation of the extension. Each thanks is logged to a database and is associated with the installation ID of the user, the line number, the line of code, and a personal note if provided. The implementation of the extension is open source and available at: [anonymized for review]. Implementation notes are in Appendix A.
@@
By providing a communication affordance within the code editor, and allowing users to express appreciation at the level of packages, modules, or functions rather than individual commits or pull requests, the extension lowers procedural effort in expressing appreciation. We chose to present the communication affordance via a persistent button on the interface to provide users with a subtle reminder of the contributors.
@@ 4.2.1. Participants
Two criteria were used to screen potential participants. We required that they primarily use VS Code as their development environment and that they anticipated writing Python/JavaScript/TypeScript code at least 3 times a week. The screening survey also included questions about users’ demographics and backgrounds. Responses are summarized in Table 1. The 18 users (12 men and 6 women) were aged 23-30. The survey asked users to report their experience levels as either Beginner, Advanced Beginner, Intermediate, Advanced, or Expert (categories present in the developer skill matrix). As Table 1 shows, users described themselves as Advanced (13/28), Intermediate (4/18), and Expert (1/18). We also asked users to report their weekly average of the amount of time they spent writing code, which ranged from 5 hours to 45 hours a week.
@@ 4.2.2. Procedure
We deployed the extension for three weeks so that users would have sufficient time to familiarize themselves with it and explore how they might interact with it in the course of their usual programming activities. To observe naturalistic use, required usage was deliberately kept minimal; however, we still encouraged (without enforcing) users to engage with the extension a meaningful amount so that they had opportunities to reveal their experiences, and so that there would be sufficient messages with which to understand contributor experiences. Each user was required to participate in a 15-minute onboarding session conducted over video call, during which we helped them install the extension and provided a brief tutorial. During this session, they were also required to complete a brief pre-study questionnaire that asked them to provide: (1) an open-ended response describing how often they typically thanked contributors of packages they used; and (2) a rating on a seven-point scale (strongly disagree to strongly agree) indicating their agreement with the statement: “There are many developers whose work I am grateful for.” These questions were intended to capture their current feelings and practices of expressing appreciation. These are presented in Table 3 (Appendix B) to provide more detail about the participant population. At the end of the onboarding session, we encouraged users to send thanks at least two times every day they found themselves coding while expressing that this would not be enforced. Users were compensated \$20 for participating in the onboarding session and completing the pre-study questionnaire. During the three weeks, for every thanks, the extension logged the installation ID [^11] For the purpose of the study, we established a link between participants and their installation IDs during the onboarding session., line of code, line number, and personal note if added. We planned to use the thanks logged over the course of the deployment to pursue our research activities with contributors. We also planned to analyze usage patterns to address RQ2. To further investigate our research questions, at the end of the three weeks, we invited all 18 users to participate in a one-on-one interview, 14 of whom agreed to participate (those who did not are indicated in Table 1). Interviews were conducted in English, remotely via Zoom, and lasted between 20 and 50 minutes, for which users were compensated \$15. Interviews were semi-structured and investigated factors that contributed to users’ feelings of appreciation and decisions to express it, the experience of selecting what to thank and what to include in personal notes, trends in their usage, and feedback on the interface and interaction. Across these topics, we asked users if there were ways in which different or new designs could have better supported their experience. All interviews were recorded and transcribed.
@@ 4.3.1. Identifying contributors:
We first identified each unique object that was thanked (the content of the line of code at which the button was clicked). For this process, we treated packages, modules, and functions as separate objects since the 20 most recent contributors to each could be different. Thanks logged next to the import of a package were treated as thanks directed at the package as a whole. Similarly, thanks logged next to imports of modules, and next to function calls, are treated as being directed at the module and function respectively. We note that each object could have been thanked multiple times (e.g. “import cv2” was thanked twice). Corresponding to each unique object, we maintained a count of the number of thanks it was associated with and all personal notes associated with that object. We identified a total of $7070$ unique objects that were thanked, most just once. We then mapped each object of appreciation to its source code on GitHub. Thanked packages were mapped to their repositories, while thanked modules and functions were mapped to their corresponding file. We used this to then find the 20 most recent contributors. For thanked packages, we identified the 20 most recent contributors to the entire repository. For thanked modules and functions, we identified the 20 most recent contributors to the corresponding file. Every commit, contains an email address of the author of the code patch, which provided us with contact information for the contributors. In determining the 20 most recent contributors, we skipped commits where the provided email address was anonymized (emails ending in “users.noreply.github.com”). If the total number of unique contributors was fewer than 20, we recorded all contributors. This process resulted in a total of $550550$ contributors to be notified. $470470$ had been thanked for $11$ object each, with the remaining being thanked for $22$ to $88$ objects each.
@@ 4.3.2. Notification
As shown in Figure 3 (B), the notification was organized so that each segment showed one object that the contributor had been thanked for (B1). Above the object, the notification showed the number of thanks corresponding to it (B3). Below the object, we included any personal notes associated with it (B4). The notification was delivered via an email (as shown in Figure 3 (A)) that provided the contributors with context about our project so that they could understand the conditions under which the users had directed these thanks and personal notes towards them. The design of the notification message was modeled after notifications in Wikipedia’s “Thanks” system (Matias et al., [n. d.]).
Figure 3. (A) The Hug Report notification was delivered via an email that provided the contributors with context about our project so that they could understand the conditions under which the users had directed these thanks and personal notes towards them. (B) The Hug Report notification, itself, was organized so that each segment (B1) showed one object that the contributor had been thanked for (B2). Above the object, the notification showed the number of thanks corresponding to it (B3). Below the object, were any personal notes associated with it (B4)